OpenStudy (anonymous):

given the annual interest rate and the compounding period, find i, the interest rate per compounding period: 5% compounded monthly

OpenStudy (anonymous):

I do not remember the formula for compounding so I will explain it. So when you are compounding each month, the principal amount on which you calculate interest is the (principal amount + interest) from previous month For first month let the Principal Amount be P and interest rate be i. So the interest = P*i/100. So the principal amount for second month = P + P*i/100 = P(1+i/100)

OpenStudy (anonymous):

So for the second month interest = P(1+i/100)*(i/100) so the amount left after adding the interest = P(1+i/100) + P(1+i/100)*(i/100) = P(1+i/100)*(1+i/100)

OpenStudy (anonymous):

= P(1+i/100)^2 So for n months we can generalize by saying the money after n months = P(1+i/100)^n So after one year (12 months) it is P(1+1/100)^12

OpenStudy (anonymous):

Annual Interest Rate = 5% so if your money if compounded annually and taking the same starting amount (P) after one year you will have P(1+5/100)

OpenStudy (anonymous):

Since in both ways the final amount should be same so, P*(1+5/100) = P*(1+i/100)^12

OpenStudy (anonymous):